12x+15y=36 and 7x-12y=187
Answer by TimothyLamb(4379)

12x+15y=36
7x-12y=187
---
put the system of linear equations into standard form
---
12x+15y=36
7x-12y=187
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 13
y = -8
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php